OpenStudy (anonymous):

g(x)= ((3x^2-2)/(2x+3))^3 Find the Derivative, please. My math book didn't include the please.

OpenStudy (amistre64):

did it include the product rule? or quotient rule?

OpenStudy (amistre64):

\[g(x)= \left(\frac{3x^2-2}{2x+3}\right)^3\]

OpenStudy (amistre64):

i see power, chain, quotient, and constant rules .... for the most part

OpenStudy (amistre64):

your math book is expecting you to apply what youve learned ....

OpenStudy (anonymous):

The function looks exactly like you wrote it and I get all jumbled up between using the different rules. I've been able to do the other problems, I think my mistake is when I try using the constant rule, but I'm not sure.

OpenStudy (amistre64):

start with the chain rule

OpenStudy (amistre64):

\[g=(\frac{p}{q})^n\] \[g'=n(\frac{p}{q})^{n-1}~~(\frac{p}{q})'\]

OpenStudy (amistre64):

the quotient rule is all thats left now\[\frac{p}{q}\to \frac{qp'-q'p}{q^2}\]

OpenStudy (anonymous):

So I use the quotient rule to find the second part that I need to put into the chain rule, so the (p/q)' is found with the quotient rule?

OpenStudy (amistre64):

yes, the quotient rule is used to play with "fractions"